# A Survey of Large Language Models in Psychotherapy: Current Landscape and Future Directions

### v1

@@ Abstract
Mental health remains a critical global challenge, with increasing demand for accessible, effective interventions. Large language models (LLMs) offer promising solutions in psychotherapy by enhancing the assessment, diagnosis, and treatment of mental health conditions through dynamic, context-aware interactions. This survey provides a comprehensive overview of the current landscape of LLM applications in psychotherapy, highlighting the roles of LLMs in symptom detection, severity estimation, cognitive assessment, and therapeutic interventions. We present a novel conceptual taxonomy to organize the psychotherapy process into three core components: assessment, diagnosis, and treatment, and examine the challenges and advancements in each area. The survey also addresses key research gaps, including linguistic biases, limited disorder coverage, and underrepresented therapeutic models. Finally, we discuss future directions to integrate LLMs into a holistic, end-to-end psychotherapy framework, addressing the evolving nature of mental health conditions and fostering more inclusive, personalized care.
@@ A Survey of Large Language Models in Psychotherapy:
Hongbin Nathanks: Equal contribution., Yining Hua11footnotemark: 1, Zimu Wang11footnotemark: 1, Tao Shen, Beibei Yu, Lilin Wang, Wei Wang, John Torous, Ling Chen Australian Artificial Intelligence Institute, University of Technology Sydney Harvard University   Xi’an Jiaotong-Liverpool University   University of Pennsylvania hongbin.na@student.uts.edu.au, yininghua@g.harvard.edu zimu.wang19@student.xjtlu.edu.cn
@@ 1 Introduction
Mental health plays an increasingly critical role in current healthcare and social well-being. The high prevalence of common psychological disorders, such as depression and anxiety, has led to a growing demand for accessible and effective psychological interventions. However, the core of psychotherapy resides in dynamic, contextual interpersonal interactions—therapists should continuously assess and adjust their intervention strategies Wampold and Imel (2015) based on the patient’s emotional fluctuations, verbal expressions, and social background, fostering a strong therapeutic alliance Stubbe (2018) to achieve symptom resilience. This deep and flexible process contrasts sharply with traditional NLP, which is typically limited to static or single-task settings.
@@
Large language models (LLMs) offer a new perspective to addressing this challenge. By leveraging their capability to model extensive context and perform multi-turn reasoning Wang et al. (2024e); Li et al. (2024b), LLMs can capture rich semantics and emotional signals in dialogues Ma et al. (2025), enabling end-to-end language understanding and generation. In assessment, LLMs can extract potential symptom cues from vague and fragmented expressions Tu et al. (2024); Qiu et al. (2024). During diagnosis, they integrate subjective and objective patient information across multiple utterances Chen et al. (2023a); Ren et al. (2024). In therapeutic interventions, they adapt conversational strategies based on patients’ real-time feedback, enabling more flexible and human-like interactions compared to traditional scripted systems Lee et al. (2024b, d). As a result, LLMs have the potential to surpass the conventional “discrete label recognition” paradigm, evolving toward a model of continuous, progressive clinical reasoning, enabling seamless connections across assessment, diagnosis, and treatment, aligning more closely with therapists’ cognitive process and interaction flow.
@@
Psychotherapy [ Assessment, color subtree=AssessmentColor, [ Symptoms [ Detection [ hee So et al. (2024); Tu et al. (2024); Souto et al. (2023); Raihan et al. (2024); Xu et al. (2024);
@@
Wang et al. (2024d) , leaf, text width=26.6em ] ] ] [ Cognition [ Maddela et al. (2023); Qi et al. (2024); Wang et al. (2023); Chen et al. (2023c); Gollapalli et al. (2023); Jiang et al. (2024a)
Lim et al. (2024) , leaf, text width=34.5em ] ] [ Behavior [ Li et al. (2024c); Hoang et al. (2024); Sun et al. (2024); Cohen et al. (2024) , leaf, text width=34.5em ] ] [ Advanced [ Yang et al. (2024b); Wang et al. (2024c); Srivastava et al. (2024) , leaf, text width=34.5em ] ] ] [ Diagnosis, color subtree=DiagnosisColor [ Static Diagnosis [ Galatzer-Levy et al. (2023); Jiang et al. (2024b); Hengle et al. (2024); Lan et al. (2024c); Kuzmin et al. (2024) , leaf, text width=34.5em ] ] [ Dynamic Diagnosis [ Chen et al. (2023a); Lan et al. (2024b); Ren et al. (2024); Lan et al. (2024a) , leaf, text width=34.5em ] ] ] [ Treatment, color subtree=TreatmentColor [ LLM as a Virtual
@@
Tool [ Welivita and Pu (2023); Sharma et al. (2023); Maddela et al. (2023); Moon and Bhattacharyya (2024); Lin et al. (2024);
Na (2024); Brown et al. (2024); Zhan et al. (2024) , leaf, text width=34.5em ] ] [ LLM as Simulated
@@
and Quality Analysis [ Lee et al. (2024e); Zhang et al. (2024a); Wang et al. (2024a); Chiu et al. (2024); Li et al. (2024a); Shapira and Alfi-Yogev (2024);
Sun et al. (2024); Cohen et al. (2024); Na et al. (2024); Nguyen et al. (2024); Zhang et al. (2024b) , leaf, text width=34.5em ] ] ] ]
@@
However, existing research on applying LLMs in this field remains somewhat disjointed. Many studies have utilized LLMs for isolated tasks, such as depression detection Yang et al. (2023); Souto et al. (2023) or diagnosis Jiang et al. (2024b), regarding them as superior feature extractors. Another research line has focused on developing mental health counseling chatbots Chen et al. (2023b); Zhang et al. (2024a); however, these systems remain limited to partial assistance due to insufficient integration with clinical workflows. In other words, although LLMs hold the potential to span the entire continuum from assessment to intervention, they remain limited by the fragmented paradigms of traditional NLP, preventing them from fully leveraging their dynamic, contextual capabilities.
To address these gaps, we introduce the first taxonomy that divides the psychotherapy process into three essential dimensions: Assessment, Diagnosis, and Treatment and provides a systematic review of the recent advancements and challenges of LLMs in each stage. We further examine the current landscape from various perspectives, including the coverage of mental disorders, diversity of linguistic resources, alignment with psychotherapy theories, and the types of techniques employed, thereby sketching the overall distribution and characteristics of existing research. Building on this foundation, we discuss key challenges for the future, including issues of technical coherence, resource and language imbalances, and the disconnect between LLM-based approaches and established psychological practices. Through this comprehensive review and framework, we aim to offer methodological insights to inform future research and facilitate the practical integration of intelligent systems across the entire psychotherapy process.
@@ 2 Conceptual Taxonomy
To establish a standardized framework for understanding psychotherapy, we propose a hierarchical taxonomy aligned with the American Psychological Association’s tripartite model of psychotherapeutic processes [^1] https://www.apa.org/topics/psychotherapy. As illustrated in Figure 1, this taxonomy organizes psychotherapy into three core components: (1) Assessment, (2) Diagnosis, and (3) Treatment, with dynamic interconnections [^2] Throughout this taxonomy, the terms Assessment, Diagnosis, and Treatment specifically refer to the three core components of psychotherapy.. Each component is detailed below.
@@ 2.4 Interrelations
The taxonomy’s components interact through three dynamic processes that define psychotherapy as a complex adaptive system:
@@ 3.1 Assessment
hee So et al. (2024) Multi-turn Dialogue Fine-Tuning MLC/IE/SUM Multiple Symptoms
@@
Souto et al. (2023) Single Post Fine-Tuning MLC/EG Depression
@@
Wang et al. (2024d) Posts From One User Zero-Shot Prompting TR Depression
@@
leverages LLMs to identify mental health conditions including depression, anxiety, PTSD, and suicidal ideation, demonstrating robust performance and multidimensional applicability across diverse scenarios. Yang et al. (2023) systematically evaluated GPT-3.5, InstructGPT3, and LLaMA models across 11 datasets, revealing that emotion-enhanced chain-of-thought prompting improves interpretability yet remains inferior to specialized supervised methods. hee So et al. (2024) achieved 70.8% zero-shot symptom retrieval accuracy in Korean psychiatric interviews using GPT-4 Turbo, while their fine-tuned GPT-3.5 attained 0.817 multi-label classification accuracy. Clinical applications show particular promise, as Tu et al. (2024) leveraged GPT-4 and Llama-2 to automate PTSD assessments through information extraction from 411 interviews, significantly enhancing diagnostic practicality.
Social media analysis benefits from approaches like Souto et al. (2023)’s interpretable depression detection framework, which demonstrated strong performance across Vicuna-13B and GPT-3.5 environments. Resource development advances include Raihan et al. (2024)’s MentalHelp dataset with 14 million instances, validated through GPT-3.5 zero-shot evaluations. For suicidal ideation monitoring, Gyanendro Singh et al. (2024) and Uluslu et al. (2024) achieved state-of-the-art evidence extraction in the CLPsych 2024 shared task through innovative prompting strategies. Open-source initiatives like MentaLLaMA by Yang et al. (2024a) and Mental-LLM by Xu et al. (2024) enable multi-symptom detection via instruction-tuned LLaMA variants, though Mohammadi et al. (2024)’s WellDunn framework reveals persistent gaps in GPT-family models’ explanation consistency.
@@
focuses on estimating the level of mental health condition intensity, particularly for depression, anxiety, and PTSD. Clinical evaluations reveal Med-PaLM 2’s zero-shot depression scoring attains clinician-level alignment on interview data (Galatzer-Levy et al., 2023), though with limited PTSD generalizability. When benchmarked against specialized Transformers on DAIC-WOZ dataset Gratch et al. (2014), ChatGPT and Llama-2 exhibit moderate efficacy (Arcan et al., 2024), suggesting domain-specific architectures retain advantages in structured assessments. Shifting attention to social media data, Aragon et al. (2024) proposed a pipeline that retrieves depression-relevant text, summarizes it according to the Beck Depression Inventory (BDI) Jackson-Koku (2016), and then utilizes LLMs to predict symptom severity, achieving performance similar to expert evaluations on certain measures. In a similar vein, Wang et al. (2024d) introduced an explainable depression detection system that leverages multiple open-source LLMs to generate BDI-based answers, reporting near state-of-the-art performance without additional training data. Cross-lingual extensions emerge through Skianis et al. (2024)’s framework enabling severity prediction across 6 languages and 2 mental conditions.
@@
centers on identifying and understanding maladaptive thinking patterns, such as cognitive distortions and early maladaptive schemas, using LLMs. Maddela et al. (2023) introduced a cognitive distortion dataset and employed a few-shot strategy with GPT-3.5 to generate, classify, and reframe them, while Qi et al. (2024) constructed two Chinese social media benchmarks for cognitive distortion detection and suicidal risk assessment, demonstrating that fine-tuned LLMs are more closely than zero-/few-shot methods to supervised baselines. In a related effort, Wang et al. (2023) released the C2D2 dataset containing 7,500 Chinese sentences with distorted thinking patterns. Expanding on detection methods, Chen et al. (2023c) proposed a Diagnosis of Thought prompting approach for GPT-4 and ChatGPT, which breaks down patient utterances into factual versus subjective content and supports the generation of interpretable diagnostic reasoning. Beyond cognitive distortions, Gollapalli et al. (2023) investigated zero-shot approaches with GPT-3.5 to identify early maladaptive schemas in mental health forums, highlighting challenges in label interpretability and prompt sensitivity. Complementarily, Jiang et al. (2024a) presented a hierarchical classification and summarization pipeline to extract cognitive pathways from Chinese social media text, underscoring GPT-4’s strong performance albeit with occasional hallucinations. Finally, Lim et al. (2024) introduced a multi-agent debate framework for cognitive distortion classification, reporting substantial gains in both accuracy and specificity by synthesizing multiple LLM opinions before forming a final verdict.
@@
highlights how user actions—or in the case of Motivational Interviewing (MI), language itself—can serve as a measurable indicator of one’s readiness for change. For instance, Li et al. (2024c) introduced the MAIMS framework, employing mental scales in a zero-shot setting to identify interpersonal risk factors on social media, thereby enhancing both interpretability and accuracy. In clinical dialogues, Hoang et al. (2024) demonstrated how LLMs can automatically detect a client’s motivational direction (e.g., change versus sustain talk) and commitment level, offering valuable insights for MI-based interventions. Extending such analyses to bilingual settings, Sun et al. (2024) proposed the BiMISC dataset and prompt strategies that enable LLMs to code MI behaviors across multiple languages with expert-level performance. Lastly, Cohen et al. (2024) presented MI-TAGS for automated annotation of global MI scores, illustrating how context-sensitive modeling can approximate human annotations in psychotherapy transcripts.
@@
research has evolved beyond foundational assessment tasks to emphasize novel methodological paradigms, bias mitigation, and domain-specific summarization frameworks. For instance, Yang et al. (2024b) introduced PsychoGAT—an interactive, game-based approach that transforms standardized psychometric instruments into engaging narrative experiences, improving psychometric reliability, construct validity, and user satisfaction when measuring constructs such as depression, cognitive distortions, and personality traits. In parallel, Wang et al. (2024c) systematically investigated potential biases in various LLMs across multiple mental health datasets, revealing that even high-performing models exhibit unfairness related to demographic factors. The authors proposed fairness-aware prompts to substantially reduce such biases without sacrificing predictive accuracy. Furthermore, Srivastava et al. (2024) presented the PIECE framework, which adopts a planning-based approach to domain-aligned counseling summarization, structuring and filtering conversation content before integrating domain knowledge.

### v2

@@ Abstract
Mental health is increasingly critical in contemporary healthcare, with psychotherapy demanding dynamic, context-sensitive interactions that traditional NLP methods struggle to capture. Large Language Models (LLMs) offer significant potential for addressing this gap due to their ability to handle extensive context and multi-turn reasoning. This review introduces a conceptual taxonomy dividing psychotherapy into interconnected stages–assessment, diagnosis, and treatment–to systematically examine LLM advancements and challenges. Our comprehensive analysis reveals imbalances in current research, such as a focus on common disorders, linguistic biases, fragmented methods, and limited theoretical integration. We identify critical challenges including capturing dynamic symptom fluctuations, overcoming linguistic and cultural biases, and ensuring diagnostic reliability. Highlighting future directions, we advocate for continuous multi-stage modeling, real-time adaptive systems grounded in psychological theory, and diversified research covering broader mental disorders and therapeutic approaches, aiming toward more holistic and clinically integrated psychotherapy LLMs systems.
@@ A Survey of Large Language Models in Psychotherapy:
[…]
@@ 1 Introduction
Mental health plays an increasingly critical role in current healthcare and social well-being. The high prevalence of common psychological disorders, such as depression and anxiety, has led to a growing demand for accessible and effective psychotherapy. The core of psychotherapy resides in dynamic, contextual interpersonal interactions—therapists should continuously assess and adjust their intervention strategies Wampold and Imel (2015) based on patients’ emotional fluctuations, verbal expressions, and social backgrounds, fostering a strong therapeutic alliance Stubbe (2018) to achieve symptom resilience. This deep and flexible process contrasts sharply with traditional NLP, which is typically limited to static or single-task settings.
@@
Large language models (LLMs) offer a new perspective to addressing this challenge. By leveraging their capability to model extensive context and perform multi-turn reasoning Wang et al. (2024f); Li et al. (2024b), LLMs can capture rich semantics and emotional signals in dialogues Ma et al. (2025), enabling end-to-end language understanding and generation Wang et al. (2024c); Qian et al. (2024). In assessment, LLMs can extract potential symptom cues from vague and fragmented expressions Tu et al. (2024); Qiu et al. (2024). During diagnosis, they integrate subjective and objective patient information across multiple utterances Chen et al. (2023a); Ren et al. (2024). In therapeutic interventions, they adapt conversational strategies based on patients’ real-time feedback, enabling more flexible and human-like interactions compared to traditional scripted systems Lee et al. (2024b, d). As a result, LLMs have the potential to surpass the conventional “discrete label recognition” paradigm, evolving toward a model of continuous, progressive clinical reasoning, enabling seamless connections across assessment, diagnosis, and treatment, aligning more closely with therapists’ cognitive process and interaction flow.
@@
Psychotherapy [ Assessment, color subtree=AssessmentColor, [ Symptoms [ Detection [ So et al. (2024); Tu et al. (2024); Bao et al. (2024); Raihan et al. (2024); Xu et al. (2024);
@@
Wang et al. (2024e) , leaf, text width=26.6em ] ] ] [ Cognition [ Maddela et al. (2023); Qi et al. (2024); Wang et al. (2023); Chen et al. (2023c); Gollapalli et al. (2023); Jiang et al. (2024a)
Lim et al. (2024) , leaf, text width=34.5em ] ] [ Behavior [ Li et al. (2024c); Hoang et al. (2024); Sun et al. (2024); Cohen et al. (2024) , leaf, text width=34.5em ] ] [ Advanced [ Yang et al. (2024b); Wang et al. (2024d); Srivastava et al. (2024) , leaf, text width=34.5em ] ] ] [ Diagnosis, color subtree=DiagnosisColor [ Static Diagnosis [ Galatzer-Levy et al. (2023); Jiang et al. (2024c); Hengle et al. (2024); Lan et al. (2024c); Kuzmin et al. (2024) , leaf, text width=34.5em ] ] [ Dynamic Diagnosis [ Chen et al. (2023a); Lan et al. (2024b); Ren et al. (2024); Lan et al. (2024a) , leaf, text width=34.5em ] ] ] [ Treatment, color subtree=TreatmentColor [ LLM as a Virtual
@@
Tool [ Welivita and Pu (2023); Sharma et al. (2023); Maddela et al. (2023); Moon and Bhattacharyya (2024); 43;
Na (2024); Brown et al. (2024); Zhan et al. , leaf, text width=34.5em ] ] [ LLM as Simulated
@@
and Quality Analysis [ Lee et al. (2024e); Zhang et al. (2024); Wang et al. (2024a); Chiu et al. (2024); Li et al. (2024a); Shapira and Alfi-Yogev (2024);
Sun et al. (2024); Cohen et al. (2024); Na et al. (2024); Nguyen et al. (2024); Zhang et al. (2025) , leaf, text width=34.5em ] ] ] ]
@@
However, existing research on applying LLMs in this field remains somewhat fragmented. Many studies have utilized LLMs for isolated tasks, such as depression detection Yang et al. (2023); Bao et al. (2024) or diagnosis Jiang et al. (2024c), regarding them as superior feature extractors. Another research line has focused on developing mental health counseling chatbots Chen et al. (2023b); Zhang et al. (2024); however, these systems remain limited to partial assistance due to insufficient integration with clinical workflows. In other words, although LLMs hold the potential to span the entire continuum from assessment to intervention, they remain limited by the fragmented paradigms of traditional NLP, preventing them from fully leveraging their dynamic, contextual capabilities.
To address these gaps, we introduce the first conceptual taxonomy that divides the psychotherapy process into three interconnected dimensions: Assessment, Diagnosis, and Treatment, and systematically review recent advancements and critical challenges of applying LLMs at each stage. We provide an extensive analysis of the current landscape from multiple perspectives, including the distribution of research across different psychotherapy stages, the coverage of mental disorders, the diversity of linguistic resources, and the incorporation of psychotherapy theories. Moreover, we critically evaluate the fragmented nature of existing approaches, highlighting the inadequacies in capturing dynamic symptom representations, the inherent limitations due to linguistic resource biases and problematic translations, and the diagnostic risks affecting clinical acceptance. Building on these findings, we outline essential future directions, emphasizing the need for continuous multi-stage modeling for coherent patient state tracking, real-time adaptability grounded explicitly in psychological theory, and a broadened scope of mental disorders and therapeutic frameworks. Through this comprehensive review, we aim to offer detailed methodological insights, guiding future research efforts and facilitating the practical, continuous, and theoretically-grounded integration of LLMs across the full spectrum of psychotherapy.
@@ 2 Conceptual Taxonomy
To establish a standardized framework for understanding psychotherapy, we propose a hierarchical taxonomy aligned with the American Psychological Association (APA)’s tripartite model of psychotherapeutic processes [^1] https://www.apa.org/topics/psychotherapy. As illustrated in Figure 1, this taxonomy organizes psychotherapy into three core components: (1) Assessment, (2) Diagnosis, and (3) Treatment, with dynamic interconnections [^2] Throughout this taxonomy, the terms Assessment, Diagnosis, and Treatment specifically refer to the three core components of psychotherapy.. Each component is detailed below.
@@ 2.4 Interrelations
The taxonomy’s components interact through three dynamic processes (see Figure 1) that define psychotherapy as a complex adaptive system:
@@
2.5 Scope of This Survey
Recent surveys at the intersection of artificial intelligence and mental health primarily cover broad NLP-driven interventions Malgaroli et al. (2023) or generic AI applications in cognitive behavioral therapy Jiang et al. (2024b), without specific emphasis on LLMs. Other reviews explicitly focusing on LLMs, such as the scoping review Hua et al. (2024, 2025) and the overview of general opportunities and risks Lawrence et al. (2024), examine general mental health rather than psychotherapy specifically. In contrast, our survey explicitly targets recent LLM applications within psychotherapy from the emergence of ChatGPT in late 2022 through October 2024, mainly including papers published in computational linguistics conferences and recent arXiv preprints. We adopt a slightly broad definition of LLMs, primarily including language models exceeding 7 billion parameters Peng et al. (2023); Zhao et al. (2025). Using the APA’s tripartite model as a foundation, we manually classify each paper according to psychotherapy-oriented components—Assessment, Diagnosis, and Treatment–clearly highlighting critical research gaps and future directions distinct from previous reviews.
@@ 3.1 Assessment
So et al. (2024) Multi-turn Dialogue Fine-Tuning MLC/IE/SUM Multiple Symptoms
@@
Bao et al. (2024) Single Post Fine-Tuning MLC/EG Depression
@@
Wang et al. (2024e) Posts From One User Zero-Shot Prompting TR Depression
@@
leverages LLMs to identify mental health conditions including depression, anxiety, PTSD, and suicidal ideation, demonstrating robust performance and multidimensional applicability across diverse scenarios. Yang et al. (2023) systematically evaluated GPT-3.5, InstructGPT3, and LLaMA models across 11 datasets, revealing that emotion-enhanced chain-of-thought prompting improves interpretability yet remains inferior to specialized supervised methods. So et al. (2024) achieved 70.8% zero-shot symptom retrieval accuracy in Korean psychiatric interviews using GPT-4 Turbo, while their fine-tuned GPT-3.5 attained 0.817 multi-label classification accuracy. Clinical applications show particular promise, as Tu et al. (2024) leveraged GPT-4 and Llama-2 to automate PTSD assessments through information extraction from 411 interviews, significantly enhancing diagnostic practicality.
Social media analysis benefits from approaches like Bao et al. (2024)’s interpretable depression detection framework, which demonstrated strong performance across Vicuna-13B and GPT-3.5 environments. Resource development advances include Raihan et al. (2024)’s MentalHelp dataset with 14 million instances, validated through GPT-3.5 zero-shot evaluations. For suicidal ideation monitoring, Gyanendro Singh et al. (2024) and Uluslu et al. (2024) achieved state-of-the-art evidence extraction in the CLPsych 2024 shared task through innovative prompting strategies. Open-source initiatives like MentaLLaMA by Yang et al. (2024a) and Mental-LLM by Xu et al. (2024) enable multi-symptom detection via instruction-tuned LLaMA variants, though Mohammadi et al. (2024)’s WellDunn framework reveals persistent gaps in GPT-family models’ explanation consistency.
@@
focuses on estimating the level of mental health condition intensity, particularly for depression, anxiety, and PTSD. Clinical evaluations reveal Med-PaLM 2’s zero-shot depression scoring attains clinician-level alignment on interview data (Galatzer-Levy et al., 2023), though with limited PTSD generalizability. When benchmarked against specialized Transformers on DAIC-WOZ dataset Gratch et al. (2014), ChatGPT and Llama-2 exhibit moderate efficacy (Arcan et al., 2024), suggesting domain-specific architectures retain advantages in structured assessments. Shifting attention to social media data, Aragon et al. (2024) proposed a pipeline that retrieves depression-relevant text, summarizes it according to the Beck Depression Inventory (BDI) Jackson-Koku (2016), and then utilizes LLMs to predict symptom severity, achieving performance similar to expert evaluations on certain measures. In a similar vein, Wang et al. (2024e) introduced an explainable depression detection system that leverages multiple open-source LLMs to generate BDI-based answers, reporting near state-of-the-art performance without additional training data. Cross-lingual extensions emerge through Skianis et al. (2024)’s framework enabling severity prediction across 6 languages and 2 mental conditions.
@@
centers on identifying and understanding maladaptive thinking patterns, such as cognitive distortions and early maladaptive schemas, using LLMs. Maddela et al. (2023) introduced a cognitive distortion dataset and employed a few-shot strategy with GPT-3.5 to generate, classify, and reframe them, while Qi et al. (2024) constructed two Chinese social media benchmarks for cognitive distortion detection and suicidal risk assessment, demonstrating that fine-tuned LLMs are more closely than zero-/few-shot methods to supervised baselines. In a related effort, Wang et al. (2023) released the C2D2 dataset containing 7,500 Chinese sentences with distorted thinking patterns. Expanding on detection methods, Chen et al. (2023c) proposed a Diagnosis of Thought (DoT) prompting approach for GPT-4 and ChatGPT, which breaks down patient utterances into factual versus subjective content and supports the generation of interpretable diagnostic reasoning. Beyond cognitive distortions, Gollapalli et al. (2023) investigated zero-shot approaches with GPT-3.5 to identify early maladaptive schemas in mental health forums, highlighting challenges in label interpretability and prompt sensitivity. Complementarily, Jiang et al. (2024a) presented a hierarchical classification and summarization pipeline to extract cognitive pathways from Chinese social media text, underscoring GPT-4’s strong performance albeit with occasional hallucinations. Finally, Lim et al. (2024) introduced a multi-agent debate framework for cognitive distortion classification, reporting substantial gains in both accuracy and specificity by synthesizing multiple LLM opinions before forming a final verdict.
@@
highlights how user actions–or in the case of Motivational Interviewing (MI), language itself–can serve as a measurable indicator of one’s readiness for change. For instance, Li et al. (2024c) introduced the MAIMS framework, employing mental scales in a zero-shot setting to identify interpersonal risk factors on social media, thereby enhancing both interpretability and accuracy. In clinical dialogues, Hoang et al. (2024) demonstrated how LLMs can automatically detect a client’s motivational direction (e.g., change versus sustain talk) and commitment level, offering valuable insights for MI-based interventions. Extending such analyses to bilingual settings, Sun et al. (2024) proposed the BiMISC dataset and prompt strategies that enable LLMs to code MI behaviors across multiple languages with expert-level performance. Lastly, Cohen et al. (2024) presented MI-TAGS for automated annotation of global MI scores, illustrating how context-sensitive modeling can approximate human annotations in psychotherapy transcripts.
@@
research has evolved beyond foundational assessment tasks to emphasize novel methodological paradigms, bias mitigation, and domain-specific summarization frameworks. For instance, Yang et al. (2024b) introduced PsychoGAT—an interactive, game-based approach that transforms standardized psychometric instruments into engaging narrative experiences, improving psychometric reliability, construct validity, and user satisfaction when measuring constructs such as depression, cognitive distortions, and personality traits. In parallel, Wang et al. (2024d) systematically investigated potential biases in various LLMs across multiple mental health datasets, revealing that even high-performing models exhibit unfairness related to demographic factors. The authors proposed fairness-aware prompts to substantially reduce such biases without sacrificing predictive accuracy. Furthermore, Srivastava et al. (2024) presented the PIECE framework, which adopts a planning-based approach to domain-aligned counseling summarization, structuring and filtering conversation content before integrating domain knowledge.